OpenStudy (anonymous):

The hypotenuse of a 30- 60- 90 triangle has length 15. What is the length of the side opposite the 60 angle?

OpenStudy (anonymous):

ratios are \(1:\sqrt{3}:2\) for short side: long side: hypotenuse if the hypotenuse is 15, then the long side is \(\frac{15}{\sqrt{3}}\)

OpenStudy (anonymous):

the length of the side opposite the 60 angle is \[\sqrt{3}/2\] times of the hypotenuse. Therefore length of the side opposite the 60 angle is \[\sqrt{3}/2\times15=15\sqrt{3}/2 units\]